OpenStudy (anonymous):

solve. Q -7 5/6 = 6 1/2

OpenStudy (anonymous):

A. - 1/4 b. 12 2/3 c. 14 1/3 d. 12 1/2

OpenStudy (joannablackwelder):

Are those mixed numbers?

OpenStudy (anonymous):

idk :(

OpenStudy (joannablackwelder):

Does it look like this: \[q-7\frac{ 5 }{ 6}=6\frac{ 1 }{ 2 }\]

OpenStudy (joannablackwelder):

Does it look like that?